# Real-Time Neural Video Compression with Unified Intra and Inter Coding

### v4

@@ Abstract
Neural video compression (NVC) technologies have advanced rapidly in recent years, yielding state-of-the-art schemes such as DCVC-RT that offer superior compression efficiency to H.266/VVC and real-time encoding/decoding capabilities. Nonetheless, existing NVC schemes have several limitations, including inefficiency in dealing with disocclusion and new content, interframe error propagation and accumulation, among others. To eliminate these limitations, we borrow the idea from classic video coding schemes, which allow intra coding within inter-coded frames. With the intra coding tool enabled, disocclusion and new content are properly handled, and interframe error propagation is naturally intercepted without the need for manual refresh mechanisms. We present an NVC framework with unified intra and inter coding, where every frame is processed by a single model that is trained to perform intra/inter coding adaptively. Moreover, we propose a simultaneous two-frame compression design to exploit interframe redundancy not only forwardly but also backwardly. Experimental results show that our scheme outperforms DCVC-RT by an average of 12.1% BD-rate reduction, delivers more stable bitrate and quality per frame, and retains real-time encoding/decoding performances. Code and models will be released.
@@ 1 Introduction
Exploiting interframe redundancy lies at the core of video compression, as efficient removal of temporal redundancy directly translates to substantial bitrate savings. For decades, classic video coding standards [10, 31, 41, 36, 3] have continuously advanced temporal redundancy mining—for example, via improved motion models [40, 42, 24] that enhance prediction accuracy. Recently, neural video compression (NVC) frameworks [12, 34, 21, 37, 35, 22, 15, 2, 44, 38, 14, 6] have further emphasized the utilization of inter-frame information to pursue higher compression efficiency. However, a critical oversight persists across most existing NVC designs: they prioritize inter-frame redundancy exploitation while neglecting the enhancement of intra-coding capabilities for scenarios where reference information is scarce or unreliable.
This limitation manifests acutely in practical scenarios such as scene changes. For instance, when the last frame of a preceding scene and the first frame of a new scene share no temporal correlation (Fig. 1), the P-frame model is forced to rely on its intrinsic intra-coding capacity. Given that state-of-the-art (SOTA) NVC schemes lack robust intra-coding support in P-frame designs, this mismatch leads to two key issues: (1) significant quality degradation and (2) severe inter-frame error propagation—both of which compromise the quality of subsequent frames.
@@
A second major challenge arises from GPU memory constraints during NVC training. The number of frames used for model training is typically insufficient to cover the full length of ultra-long sequences encountered in real-world testing. For such extended sequences, the heavy reliance of NVC schemes on previous frame features exacerbates error accumulation in reference signals. To mitigate this, periodic feature refresh techniques [22] have been adopted in recent SOTA schemes [15, 2, 38, 14]: rich feature information is reconstructed into three-channel pixel images via a recovery network, and these images are fed back as new reference features through an adaptor to intercept error accumulation. While this approach effectively blocks prior-frame errors, it suffers from two critical drawbacks: (1) it discards valuable inter-frame information (e.g., long-term motion cues or occluded object details) alongside errors, and (2) it induces sharp bitrate surges at refresh points—posing risks of network congestion and hindering practical deployment. As shown in Fig. 1, the P-frames at refresh points exhibit bitrate patterns similar to intra-coded frames, yet existing models cannot autonomously correct propagation errors via adaptive intra-coding (due to weak P-frame intra capabilities), leaving them dependent on this inflexible refresh mechanism.
@@
To address these core limitations—insufficient P-frame intra-coding, unmanaged error propagation, and disruptive manual refresh—we propose a unified NVC framework that integrates intra- and inter-frame coding capabilities into a single model. Through training, we enable our model to adaptively balance between intra- and inter-frame coding. When reference information is accurate and abundant, the model prioritizes inter-frame prediction to maximize redundancy reduction; when references are error-prone or insufficient, it adaptively invokes intra-coding to enhance current-frame quality. This design naturally handles scene transitions, maintains stable bitrate and quality without manual refresh, and inherently mitigates inter-frame error propagation.
@@
Balancing low bitrate, high visual quality, and real-time inference performance remains a critical and non-trivial challenge in reference-scarce scenarios (e.g., scene changes). Despite advancements in low-complexity real-time (RT) neural video compression (NVC) frameworks [14], SOTA solutions still rely on computationally expensive standalone I-frame models to address these reference-scarce cases. Integrating such heavyweight intra-coding complexity directly into inter-frame coding pipelines would inevitably degrade inference speed—a key constraint for practical low-latency applications (e.g., real-time video streaming). To resolve this fundamental trade-off, we propose a simultaneous two-frame compression technique: by jointly encoding two consecutive frames, this approach leverages rich backward reference information from the subsequent frame, while incurring merely one frame latency. Moreover, the collaborative modeling of inter-frame dependencies enables the extraction of more comprehensive temporal cues (e.g., fine-grained motion trends and cross-frame object correlations) that are inaccessible in single-frame coding.
@@
  • We unify intra- and inter-frame coding into a single model, eliminating the need for a separate intra-coding model. This design enhances the NVC’s ability to handle scene changes and reduces overall model parameter count.
@@ 2 Related Work
Classic video coding standards (e.g., H.264/AVC, H.265/HEVC, H.266/VVC) [10, 31, 41, 36, 3] have long relied on inter-frame prediction to reduce temporal redundancy, where the current frame is predicted using reference frames via motion estimation and compensation. A critical design in these standards is the integration of intra prediction within inter-frame coding: when inter prediction fails to yield satisfactory performance (e.g., for disoccluded regions, newly appearing content, or areas with complex motion where motion vectors cannot capture accurate dependencies), local blocks are switched to intra coding mode. This strategy avoids excessive residual distortion caused by poor inter prediction, balancing compression efficiency and reconstruction quality—an essential heuristic for handling scene inhomogeneities that remains valuable for modern NVC schemes.
@@
Recent advances in NVC have demonstrated great potential in video compression. Existing schemes can be broadly categorized into two main types: residual coding and conditional coding. Residual coding [7, 26, 33, 32, 28, 29, 12, 13, 1] primarily reduces redundancy by encoding the difference between predicted frames and the current frame. Conditional coding [17, 18, 19, 34, 11, 20, 21, 22, 15, 2, 44, 38, 14] focuses on exploiting contextual correlations to improve coding performance. Some other emerging frameworks based on neural video representations [4, 5, 45, 23, 16, 9, 27] are also rapidly developing. Given that conditional coding can learn complex motion patterns and has entropy lower than or equal to that of residual coding—and with the advent of DCVC RT [14]—real-time conditional coding is increasingly approaching practical applicability. Our model adopts a real-time conditional coding design and introduces improvements in reducing inter-frame error propagation.
@@
Conditional coding leverages contextual information for prediction, where the prediction performance depends on the richness and accuracy of the context. Most existing NVC schemes enhance the context by efficiently utilizing temporal references [15, 2, 44, 38, 14]. For instance, DCVC-FM [22] proposes training with longer sequences and a long-sequence refresh mechanism to maintain the accuracy of the temporal context. ECVC [15] utilizes non-local temporal context to enrich the reference information for the current frame. SEVC [2] preserves richer contextual information through multi-resolution temporal contexts. However, these methods rely on optical flow networks, which often entail high computational complexity and are unsuitable for real-time applications. In contrast, DCVC-RT [14] explores temporal context correlations via implicit context alignment, eliminating the need to encode additional motion information. This approach significantly reduces computational complexity, making the practical deployment of NVC feasible. Nevertheless, DCVC-RT still faces unresolved critical challenges—including inefficiency in handling disocclusion and new content, as well as inter-frame error propagation and accumulation—which continue to restrict its further practical deployment. This technical gap serves as the key motivation for the core innovation of our work.
@@
Figure 3: Detailed architecture of our proposed UI2C (unified intra and inter coding) scheme. $(E_{1},E_{2},E_{3})$, $(D_{1},D_{2})$, $(R_{1},R_{2})$, $(FE_{1},FE_{1})$ represent the components of the encoder, decoder, reconstruction generator, and feature extractor, respectively (detailed configurations are provided in the supplementary material). The reference feature $F_{(t-2,t-1)}$ is used to obtain the contextual feature $C^{e}_{(t,t+1)}$ and $C_{(t,t+1)}$. The quantization parameters $qp_{1}$ and $qp_{2}$ are used to select the respective quantization tables. $\times$ stands for element-wise multiply.
@@ 3.1 Overview
Our proposed scheme, namely UI2C (Unified Intra and Inter Coding), builds on the real-time neural codec DCVC-RT [14], with core enhancements to address prior limitations. As shown in Fig. 2, UI2C removes dedicated I-frame models by integrating unified intra-inter coding into a single spatio-temporal network (Section 3.2). To encode $x_{t}$, when $t$ is even ($t=0,1,2,\dots$), 1-frame latency is introduced to wait for $x_{t+1}$ for simultaneous two-frame compression. These two frames are concatenated and fed to a shared encoder-decoder, exploiting forward (prior decoded frames) and backward ($x_{t+1}$ for $x_{t}$) temporal redundancies [14]. The decoder reconstructs both frames from one piece of code, retaining their features in the reference buffer for subsequent coding (Section 3.3).
To optimize rate-distortion (RD) performance across the two frames, a two-frame quantization table strategy allocates bitrates by frame indexes (Section 3.4). For training, a hybrid reference scheme enhances adaptive coding: initial batch frames randomly use blank references (simulating intra-dominant scenarios) or noise-perturbed prior frames (mimicking error-prone inter references), enabling UI2C to switch modes based on reference reliability (Section 3.5).
@@ 3.2 Unified intra- and inter-frame coding
Moreover, since the number of available P-frames for training is inherently limited, existing models are unable to automatically handle extremely long sequences without external manual intervention, such as inserting I-frames or applying periodic refresh mechanisms. Invoking the I-frame model truncates all potential errors through intra-coding, which discards all potentially useful references. This leads to the RD performance of IP32 (Intra Period=32 frames) configurations consistently underperforming compared to IP-1 (One Intra frame). The manual refresh method [22] converts accumulated features back to the pixel-domain image, which is then processed by an adaptor to generate new features. This approach preserves partial information (only spatial information from the previous frame, while all motion information is discarded) and simultaneously eliminates errors. However, this forces the subsequent first P-frame to encode with limited information, making the model treat it as the first P-frame after an I-frame, thereby initiating a new coding cycle. This lack of information compels the P-frame model to resort to more intra-coding. Yet, the current P-frame model exhibits weak intra-coding capabilities and struggles to effectively balance intra- and inter-coding, resulting in excessively high bitrates for that frame.
Based on this observation, we argue that a separate I-frame model is unnecessary. Instead, a single unified model can handle both intra- and inter-coding scenarios. During inference, for the first frame, a blank frame can be fed through an adaptor to generate reference features, enabling the direct use of the model’s intra-coding capability. As illustrated in Fig. 2, our model processes the first frame by transforming a blank frame into features via an adaptor, which serves as reference input, thus leveraging the model’s inherent intra-coding ability. For subsequent frames, the same model is reused, but now with informative reference features, allowing it to utilize its inter-coding capabilities predominantly.
@@ 3.3 Simultaneous two-frame compression
For reference-scarce cases (e.g., first frame or scene transitions), backward references from $x_{t+1}$ compensate for the absence of prior frame information, avoiding quality degradation caused by weak intra-coding under constrained complexity. For inter-coding, bidirectional cues enable more accurate modeling of occluded regions (e.g., reappearing objects) and provide error calibration for noisy propagated features.
Notably, consecutive frames in natural video sequences exhibit substantial temporal redundancy with high similarity. After 8× joint downsampling, trivial high-frequency variations between the two frames are suppressed, further enhancing their feature-level consistency to enable efficient joint encoding. As illustrated in Fig. 3, building on DCVC-RT’s efficiency-driven design [14], we concatenate $x_{t}$ and $x_{t+1}$ along the channel dimension, apply the aforementioned joint downsampling, and feed the fused feature into the shared encoder-decoder. This single-stream pipeline leverages the implicit modeling capability of neural networks for mutual reference between the two frames, generating only one compact bitstream while preserving real-time inference speed. At the decoder side, both frames are reconstructed synchronously; the fused features generated here are stored in the reference buffer, providing richer contextual information for the coding of subsequent frames and ultimately forming a performance-enhancing loop.
@@ 3.4 Two-frame quantization
To regulate the quality control between the two frames during encoding, we adopt a two-frame quantization approach. As is shown in Fig. 3, for the two input frames, each frame is assigned a quality parameter (qp) queried based on its frame index, resulting in two distinct qp values. For these two qp values, corresponding quantization coefficients are obtained by querying different quantization tables. The quantization coefficients of both frames are then concatenated and multiplied by the features at corresponding positions to achieve quality control. Moreover, we assign a higher qp to the latter frame of the two frames, so that subsequent frames have a better reference.

### v5

@@ Abstract
Neural video compression (NVC) technologies have advanced rapidly in recent years, yielding state-of-the-art schemes such as DCVC-RT that offer superior compression efficiency to H.266/VVC and real-time encoding/decoding capabilities. Nonetheless, existing NVC schemes have several limitations, including inefficiency in dealing with disocclusion and new content, interframe error propagation and accumulation, among others. To eliminate these limitations, we borrow the idea from classic video coding schemes, which allow intra coding within inter-coded frames. With the intra coding tool enabled, disocclusion and new content are properly handled, and interframe error propagation is naturally intercepted without the need for manual refresh mechanisms. We present an NVC framework with unified intra and inter coding, where every frame is processed by a single model that is trained to perform intra/inter coding adaptively. Moreover, we propose a simultaneous two-frame compression design to exploit interframe redundancy not only forwardly but also backwardly. Experimental results show that our scheme outperforms DCVC-RT by an average of 12.1% BD-rate reduction, delivers more stable bitrate and quality per frame, and retains real-time encoding/decoding performances. The codes are available at https://github.com/ihuixiang/UIIC.
@@ 1 Introduction
Exploiting interframe redundancy lies at the core of video compression, as efficient removal of temporal redundancy directly translates to substantial bitrate savings. For decades, classic video coding standards [10, 31, 42, 36, 3] have continuously advanced temporal redundancy mining via improved motion models [40, 41, 23] that enhance prediction accuracy. Recently, neural video compression (NVC) frameworks [13, 34, 21, 38, 35, 22, 15, 2, 44, 37, 14, 6] have further emphasized the utilization of inter-frame information to pursue higher compression efficiency. However, a critical oversight persists across most existing NVC designs: they prioritize exploiting inter-frame redundancy while neglecting to enhance intra-coding capabilities in scenarios where reference information is scarce or unreliable.
This limitation manifests acutely in practical scenarios such as scene changes. For instance, when the last frame of a preceding scene and the first frame of a new scene share no temporal correlation (Figure 1), the P-frame model is forced to rely on its intrinsic intra-coding capacity. Given that state-of-the-art (SOTA) NVC schemes lack robust intra-coding support in P-frame designs, this mismatch leads to two key issues: (1) significant quality degradation and (2) severe inter-frame error propagation—both of which compromise the quality of subsequent frames.
@@
A second major challenge arises from GPU memory constraints during NVC training. The number of frames used for model training is typically insufficient to cover the full length of ultra-long sequences encountered in real-world testing. For such extended sequences, the heavy reliance of NVC schemes on previous frame features exacerbates error accumulation in reference signals. To mitigate this, periodic feature refresh techniques [22] have been adopted in recent SOTA schemes [15, 2, 37, 14]: rich feature information is reconstructed into a three-channel pixel image via the recovery network, and this image is fed back as new reference features through an adaptor to intercept error accumulation. While this approach effectively blocks prior-frame errors, it suffers from two critical drawbacks: (1) it discards valuable inter-frame information (e.g., long-term temporal cues or occluded object details) alongside errors, and (2) it induces sharp bitrate surges at refresh points, posing risks of network congestion and hindering practical deployment. As shown in Figure 1, the P-frames at refresh points exhibit bitrate patterns similar to intra-coded frames, yet existing models cannot autonomously correct propagation errors via adaptive intra-coding (due to weak P-frame intra capabilities), leaving them dependent on this inflexible refresh mechanism.
@@
To address these core limitations, we propose a unified NVC framework that integrates intra- and inter-frame coding capabilities into a single model. Through training, we enable our model to adaptively balance between intra- and inter-frame coding. When reference information is accurate and abundant, the model prioritizes inter-frame prediction to maximize redundancy reduction; when references are error-prone or insufficient, it adaptively invokes intra-coding to enhance current-frame quality. This design naturally handles scene transitions, maintains stable bitrate and quality without manual refresh, and inherently mitigates inter-frame error propagation.
@@
Balancing low bitrate, high visual quality, and real-time inference performance remains a critical and non-trivial challenge in reference-scarce scenarios (e.g., scene changes). Despite advancements in low-complexity real-time (RT) neural video compression (NVC) frameworks [14], SOTA solutions still rely on computationally expensive standalone I-frame models to address these reference-scarce cases. Integrating such heavyweight intra-coding complexity directly into inter-frame coding pipelines would inevitably degrade inference speed, which is a key constraint for practical low-latency applications (e.g., real-time video streaming). To resolve this fundamental trade-off, we propose a simultaneous two-frame compression technique: by jointly encoding two consecutive frames, this approach leverages rich backward reference information from the subsequent frame, while incurring merely one frame latency. Moreover, the collaborative modeling of inter-frame dependencies enables the extraction of more comprehensive temporal cues (e.g., fine-grained temporal trends and cross-frame object correlations) that are inaccessible in single-frame coding.
@@
  • We unify intra- and inter-frame coding into a single model, eliminating the need for a separate intra-coding model. This design enhances the NVC’s ability to handle scene changes and reduces model parameter count.
@@ 2 Related Work
Classic video coding standards (e.g., H.264/AVC, H.265/HEVC, H.266/VVC) [10, 31, 42, 36, 3] have long relied on inter-frame prediction to reduce temporal redundancy, where the current frame is predicted using reference frames via motion estimation and compensation. A critical design in these standards is the integration of intra prediction within inter-frame coding: when inter prediction fails to yield satisfactory performance (e.g., for disoccluded regions, newly appearing content, or areas with complex motion where motion vectors cannot capture accurate dependencies), local blocks are switched to intra coding mode. This strategy avoids excessive residual distortion caused by poor inter prediction, balancing compression efficiency and reconstruction quality. The same principle retains significant value for modern neural video coding (NVC) frameworks, where addressing scene inhomogeneity and inter-frame error propagation remains a fundamental challenge.
@@
Recent advances in NVC have demonstrated great potential in video compression. Existing schemes can be broadly categorized into two main types: residual coding and conditional coding. Residual coding [7, 26, 32, 33, 28, 29, 13, 12, 1] primarily reduces redundancy by encoding the difference between predicted frames and the current frame. Conditional coding [17, 18, 19, 34, 11, 20, 21, 22, 15, 2, 44, 37, 14] focuses on exploiting contextual correlations to improve coding performance. Some other emerging frameworks based on neural video representations [5, 4, 45, 24, 16, 9, 27] are also rapidly developing. Given that conditional coding can learn complex temporal patterns and has entropy lower than or equal to that of residual coding, real-time conditional coding is increasingly approaching practical applicability with the advent of DCVC RT [14]. Our model adopts a real-time conditional coding design and introduces improvements in reducing inter-frame error propagation.
@@
Conditional coding leverages contextual information for prediction, where the prediction performance depends on the richness and accuracy of the context. Most existing NVC schemes enhance the context by efficiently utilizing temporal references [15, 2, 44, 37, 14]. For instance, DCVC-FM [22] proposes training with longer sequences and a long-sequence refresh mechanism to maintain the accuracy of the temporal context. ECVC [15] utilizes non-local temporal context to enrich the reference information for the current frame. SEVC [2] preserves richer contextual information through multi-resolution temporal contexts. However, these methods rely on optical flow networks, which often entail high computational complexity and are unsuitable for real-time applications. In contrast, DCVC-RT [14] explores temporal context correlations via implicit context alignment, eliminating the need to encode additional motion information. This approach significantly reduces computational complexity, making the practical deployment of NVC feasible. Nevertheless, DCVC-RT still faces unresolved critical challenges, including inefficiency in handling disocclusion and new content, as well as inter-frame error propagation and accumulation, which continue to restrict its further practical deployment. This technical gap serves as the key motivation for the core innovation of our work.
@@
Figure 3: Detailed architecture of our proposed UI2C (unified intra and inter coding) scheme. $(E_{1},E_{2},E_{3})$, $(D_{1},D_{2},D_{3})$, $(R_{1},R_{2})$, $(FE_{1},FE_{1})$ represent the components of the encoder, decoder, reconstruction generator, and feature extractor, respectively (detailed configurations are provided in the supplementary material). The reference feature $F_{(t-2,t-1)}$ is used to obtain the contextual feature $C^{e}_{(t,t+1)}$ and $C_{(t,t+1)}$. The quantization parameters $qp_{1}$ and $qp_{2}$ are used to select the respective quantization tables. $\times$ stands for element-wise multiply.
@@ 3.1 Overview
Our proposed model, namely UI2C (Unified Intra and Inter Coding), builds on the real-time neural codec DCVC-RT [14], with core enhancements to address prior limitations. As shown in Figure 2, UI2C removes dedicated I-frame models by integrating unified intra-inter coding into a single spatio-temporal network (Section 3.2). To encode $x_{t}$, when $t$ is even ($t=0,2,\dots$), 1-frame latency is introduced to wait for $x_{t+1}$ for simultaneous two-frame compression. These two frames are concatenated and fed to a shared encoder-decoder, exploiting forward (prior decoded frames) and backward ($x_{t+1}$ for $x_{t}$) temporal redundancies [14]. The decoder reconstructs both frames from one piece of code, retaining their features in the reference buffer for subsequent coding (Section 3.3).
To optimize rate-distortion (RD) performance across the two frames, a two-frame quantization table strategy allocates bitrates by frame indexes (Section 3.4). For training, a hybrid reference scheme enhances adaptive coding: initial batch frames randomly use blank references (simulating intra-dominant scenarios) or noise-perturbed prior features (mimicking error-prone inter references), enabling UI2C to switch modes based on reference reliability (Section 3.5).
@@ 3.2 Unified intra- and inter-frame coding
Moreover, since the number of available P-frames for training is inherently limited, existing models are unable to automatically handle extremely long sequences without external manual intervention, such as inserting I-frames or applying periodic refresh mechanisms. Invoking the I-frame model truncates all potential errors through intra-coding, which discards all potentially useful references. This leads to the RD performance of IP32 (Intra Period=32 frames) configurations consistently underperforming compared to IP-1 (One Intra frame). The manual refresh method [22] converts accumulated features back to the pixel-domain image, which is then processed by an adaptor to generate new features. This approach preserves partial information (only spatial information from the previous frame, while all temporal information is discarded) and simultaneously eliminates errors. However, this forces the subsequent first P-frame to encode with limited information, making the model treat it as the first P-frame after an I-frame, thereby initiating a new coding cycle. This lack of information compels the P-frame model to resort to more intra-coding. Yet, the current P-frame model exhibits weak intra-coding capabilities and struggles to effectively balance intra- and inter-coding, resulting in excessively high bitrates for that frame.
Based on this observation, we argue that a dedicated I-frame model is unnecessary. Instead, a single unified model can effectively handle both intra- and inter-coding scenarios. During inference, for the first frame, we feed a blank frame through an adaptor to generate reference features, enabling the model to directly use its inherent intra-coding capability. As illustrated in Figure 2, the model processes the first frame by converting the blank input into reference features via the adaptor, thereby activating its intra-coding behavior. For subsequent frames, the same model is reused with informative reference features, allowing it to primarily exploit its inter-coding capabilities.
@@ 3.3 Simultaneous two-frame compression
In reference-scarce cases (e.g., the first frame or scene transitions), backward references from $x_{t+1}$ compensate for the lack of preceding frame information, alleviating the quality degradation that would otherwise arise from weak intra-coding under constrained complexity. For inter-coding, the bidirectional cues enable more accurate modeling of occluded regions and offer error calibration for noisy or imperfectly propagated features.
Notably, consecutive frames in natural video sequences exhibit substantial temporal redundancy with high similarity. After 8× joint downsampling, trivial high-frequency variations between the two frames are suppressed, further enhancing their feature-level consistency to enable efficient joint encoding. As illustrated in Figure 3, building on DCVC-RT’s efficiency-driven design [14], we concatenate $x_{t}$ and $x_{t+1}$ along the channel dimension, apply the aforementioned joint downsampling, and feed the fused feature into the shared encoder-decoder. This single-stream pipeline leverages the implicit modeling capability of neural networks for mutual reference between the two frames, generating only one compact bitstream while preserving real-time inference speed. At the decoder side, both frames are reconstructed synchronously; the fused features generated here are stored in the reference buffer, providing richer contextual information for the coding of subsequent frames and ultimately forming a performance-enhancing loop.
@@ 3.4 Two-frame quantization
To regulate the quality control between the two frames during encoding, we adopt a two-frame quantization approach. As is shown in Figure 3, for the two input frames, each frame is assigned a quality parameter (qp) queried based on its frame index, resulting in two distinct qp values. For these two qp values, corresponding quantization coefficients are obtained by querying different quantization tables. The quantization coefficients of both frames are then concatenated and multiplied by the features at corresponding positions to achieve quality control. Moreover, we assign a higher qp to the latter frame of the two frames, so that subsequent frames have a better reference.